Find an expression of the function whose graph is the bottom half of the given parabola

y2-x-4y+3=0


1
Expert's answer
2021-10-27T12:23:38-0400

"\displaystyle
y^2 - x - 4y + 3 = 0\\

y^2 - 4y + 3 = x\\

y^2 - 4y = x - 3\\

y^2 - 4y + 4 = x - 3 + 4 = x + 1\\

(y - 2)^2 = x + 1\\

y = \pm\sqrt{x + 1} + 2\\

\textup{The bottom half of the parabola is}\\
y = -\sqrt{x + 1} + 2\\

\textup{Hence,}\\
\textup{domain}(y) = \{x \in \mathbb{R}: x \geq -1\}\\

\forall x \geq -1,\,\,\, -\sqrt{x + 1} \leq 0, \\

\textup{and so}\,\,\, 2 - \sqrt{x + 1} \leq 2\\

\textup{Thus,}\,\,\, y \leq 2.\,\,\,\forall x \geq 1.\\
\textup{Hence,}\\
\textup{range}(y) = \{y \in \mathbb{R}: y \leq 2\}\\"